# Complementary Time–Space Tradeoff for Self-Stabilizing Leader Election: Polynomial States Meet Sublinear Time

### v1

@@ 3.2.1 Loosely-stabilizing Leader Election
Several components of $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ assume a unique leader exists; to satisfy this, we run in parallel the LS-LE protocol of Sudo et al. [28], denoted by $\mathcal{P}_{{\textsc{LSLE}}}(\tau)$ here (line 1 in Algorithm 1). The following lemma allows us to assume, throughout the remainder of Section 3, that there is exactly one agent $a_{L}$ with $a_{L}.\mathtt{leader}=1$ and every other agent $a\in A\setminus\{a_{L}\}$ has $a.\mathtt{leader}=0$, where $\mathtt{leader}\in\{0,1\}$ is the leader-indicator variable.
Lemma 1 (immediate consequence of [28]).
@@
3.2.2 Epidemic Protocol
The one-way epidemic (or simply epidemic) protocol [7], frequently used in population protocols, propagates the maximum value of any given variable $\mathtt{var}$ to all agents. It consists of a single transition rule
where $a_{\mathrm{ini}}$ and $a_{\mathrm{res}}$ denote the initiator and responder. Angluin et al. [7] proved that this protocol stabilizes within $O(n\log n)$ interactions (i.e., $O(\log n)$ time) w.h.p.
Lemma 2 ([7]).
If the epidemic protocol for variable $\mathtt{var}$ starts from a configuration $C$ with $M=\max_{a\in A}C(a).\mathtt{var}$, then $\min_{a\in A}a.\mathtt{var}=M$ within $O(n\log n)$ interactions w.h.p.
3.2.3 Phase Clock with a Single Leader
Many variants of the phase clock have been used to loosely synchronize the population in protocol design. We use the simplest variant: the single-leader phase clock of Angluin et al. [7]. This protocol assumes a fixed unique leader. Each agent maintains a single variable $\mathtt{clock}\in[0,c_{M}\cdot T]$, where $c_{M}=\Theta(1)$ and $T=\operatorname{poly}(n)$ are protocol parameters. The initial value of $\mathtt{clock}$ is set to 00 for all agents, including the leader. [^5] Originally in [7], the leader’s clock starts at 00 and all others at $-1$. We instead initialize $\mathtt{clock}$ to 00 for every agent to simplify the protocol description; this change does not affect the correctness of Lemma 3. When an initiator $a_{\mathrm{ini}}$ and responder $a_{\mathrm{res}}$ interacts, $a_{\mathrm{ini}}$ does nothing and $a_{\mathrm{res}}$ updates its clock as follows:
@@ 3.2.3 Phase Clock with a Single Leader
We partition the interval $[-1,c_{M}\cdot T]$ into phases of uniform length $c_{M}=\Theta(1)$. Specifically, for each $i\in[-1,T]$, we say agent $a$ is in phase $i$ if $c_{M}\cdot i\leq a.\mathtt{clock}<c_{M}\cdot(i+1)$. For each $i\in[0,T]$, let $\mathcal{C}_{\mathrm{phase}}(i)$ denote the set of configurations in which all agents are in phase $i$.
Lemma 3 (Immediate consequence of Lemma 1 and Theorem 1 in [7]).
@@
                 $\mathtt{target}\in[1,n]\cup\{\bot\}$ $\mathtt{det}\in\{0,1\}$,
@@
[…]
@@
4       $a_{\mathrm{res}}.\mathtt{mode}\leftarrow a_{\mathrm{ini}}.\mathtt{mode}\leftarrow\bot$$a_{\mathrm{res}}.\mathtt{det}\leftarrow a_{\mathrm{ini}}.\mathtt{det}\leftarrow 0$
5
@@
[…]
@@
[…]
@@
17
Algorithm 1 $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ at an interaction where initiator $a_{\mathrm{ini}}$ and responder $a_{\mathrm{res}}$ meets.
3.3 Cyclic Execution of Sub-protocols
Using the phase clock [7] described in Section 3.2.3, $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ invokes the sub-protocols FindTarget, ${\textsc{Detect}}(r,\rho)$, and Rank in phases $[T_{1},T_{2}-2]$, $[T_{2},T_{3}-2]$, and $[T_{3},T_{4}-2]$, respectively (line 17), where
@@ 3.3 Cyclic Execution of Sub-protocols
and $c_{T}=\Theta(1)$ is a sufficiently large constant to ensure the correctness of each component. If an agent enters phase $T_{4}$, it raises a reset flag (lines 11–12), returning all agents to the initial phase 00 by the epidemic protocol (lines 2–6). Since each phase completes in $\Theta(n\log n)$ interactions w.h.p. by Lemma 3, one cycle of $T_{4}$ phases requires only $\Theta(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p.
A variable $\mathtt{mode}\in\{\bot,\mathsf{F},\mathsf{D},\mathsf{R}\}$ plays a key role in self-stabilization. The modes $\mathsf{F}$, $\mathsf{D}$, and $\mathsf{R}$ correspond to the sub-protocols FindTarget, Detect, and Rank, respectively. When an agent switches its $\mathtt{mode}$ from $x$ to a different mode $y\neq\bot$, it runs the initialization procedure for component $y$. The initialization procedures for Detect and Rank are given in the preambles of Algorithms 2 and 3, respectively, while the procedure for FindTarget is described in Section 3.6. Note that ${\textsc{Detect}}(r,\rho)$ requires the target rank $r$; upon entering mode $\mathsf{D}$, each agent $a$ adopts $a.\mathtt{target}$ for $r$. Mode switching occurs in phases $T_{1}-1$, $T_{2}-1$, and $T_{3}-1$, i.e., immediately before each component starts (line 15). An agent $a$ enters mode $\mathsf{R}$ only if $a.\mathtt{det}=1$, i.e., its rank-collision-detected flag is raised, so rank reassignment cannot begin until ${\textsc{Detect}}(r,\rho)$ detects a collision. Each component executes only when both interacting agents share the same mode; additionally, during mode $\mathsf{D}$ (collision detection), they must have the same target rank (lines 16–17).
The first part of the main function (Lines 2–12) resets all agents to phase 0 within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p., regardless of the initial configuration (Lemma 4). By Lemma 9, this ensures the population completes one full cycle of the three components in the next $\Theta(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p. This reset mechanism relies on two timers, $\mathtt{reset}\in[0,c_{R}\lg n]$ and $\mathtt{delay}\in[0,c_{D}\cdot\lceil\lg n\rceil]$, where a constant $c_{D}=\Theta(1)$ is chosen sufficiently larger than $c_{R}=\Theta(1)$. We omit the update rules for $\mathtt{reset}$ and $\mathtt{delay}$ (as specified inlines 2–12) here, as they follow standard techniques [14, 28] and closely mirror the PropagateReset sub-protocol of Burman et al. [14].
Note 1.
When an agent observes $\mathtt{reset}>0$, it resets $\mathtt{mode}=\bot$ and $\mathtt{det}=0$ (line 6). These resets are not required for correctness but simplify the subsequent proof.
@@
Let $\mathcal{I}$ be the set of configurations in which $a.\mathtt{clock}=a.\mathtt{reset}=a.\mathtt{delay}=\mathtt{det}=0$ and $a.\mathtt{mode}=\bot$ hold for all $a\in A$.
@@
Proof Sketch.
Follows from Lemma 1 of [3], Lemma 5 of [34], and Lemma 3.
@@
For each $X\in\{\mathsf{F},\mathsf{D},\mathsf{R}\}$, let $\mathcal{C}_{\mathrm{init}}(X)$ be the set of configurations of $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ in which every agent
@@
  2. 2.
    has all variables of the corresponding component correctly initialized.

### v2

@@ 5.1 Loosely-stabilizing Leader Election
Several components of $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ assume a unique leader exists; to satisfy this, we run in parallel the LS-LE protocol of [SEIM21], denoted by $\mathcal{P}_{{\textsc{LSLE}}}$ here (line 8 in Algorithm 2).
The following lemma allows us to assume, throughout the rest of this paper, except for the proof of Theorem 1 in Section 9, that there is exactly one agent $a_{L}$ with $a_{L}.\mathtt{leader}=1$, and every other agent $a\in A\setminus\{a_{L}\}$ has $a.\mathtt{leader}=0$, where $\mathtt{leader}\in\{0,1\}$ is the leader-indicator variable. This assumption will be removed in the proof of the main theorem using this lemma.
Lemma 5 (immediate consequence of [SEIM21]).
@@
Note 1.
Readers who are already familiar with population protocols may skip Section 5.2.
5.2 Phase Clock with a Single Leader
Many variants of the phase clock have been used to loosely synchronize the population in protocol design. We use the simplest variant: the single-leader phase clock of [AAE08]. This protocol assumes a fixed unique leader. Each agent maintains a single variable $\mathtt{clock}\in[0,c_{M}\cdot T]$, where $c_{M}=\Theta(1)$ and $T=\operatorname{poly}(n)$ are protocol parameters. The initial value of $\mathtt{clock}$ is set to 0 for all agents, including the leader. [^5] Originally in [AAE08], the leader’s clock starts at 0 and all others at 1-1. We instead initialize $\mathtt{clock}$ to 0 for every agent to simplify the protocol description; this change does not affect the correctness of Lemma 6. When an initiator $a_{\mathrm{ini}}$ and responder $a_{\mathrm{res}}$ interacts, $a_{\mathrm{ini}}$ does nothing and $a_{\mathrm{res}}$ updates its clock as follows:
@@ 5.2 Phase Clock with a Single Leader
We partition the interval $1,c_{M}\cdot T]$ into phases of uniform length $c_{M}=\Theta(1)$. Specifically, for each $i\in[-1,T]$, we say agent $a$ is in phase $i$ if $c_{M}\cdot i\leq a.\mathtt{clock}<c_{M}\cdot(i+1)$. For each $i\in[0,T]$, let $\mathcal{C}_{\mathrm{phase}}(i)$ denote the set of configurations in which all agents are in phase $i$.
Lemma 6 (Immediate consequence of Lemma 1 and Theorem 1 in [AAE08]).
@@
Note that the $\mathtt{clock}$ variable is monotonically non-decreasing in this protocol. As a result, in Lemma 6, the population enters $\mathcal{C}_{\mathrm{phase}}(1),\mathcal{C}_{\mathrm{phase}}(2),\dots,\mathcal{C}_{\mathrm{phase}}(T)$ in this order w.h.p.
@@
       $\mathtt{target}\in[1,n]$, $\mathtt{det}\in\{0,1\}$,
@@
Notation: $\mathit{Phase}(a)=\lfloor a.\mathtt{clock}/c_{M}\rfloor$
@@
14 $a_{\mathrm{res}}.\mathtt{mode}\leftarrow a_{\mathrm{ini}}.\mathtt{mode}\leftarrow\bot$; $a_{\mathrm{res}}.\mathtt{det}\leftarrow a_{\mathrm{ini}}.\mathtt{det}\leftarrow 0$
15
@@
  // Results in $\mathit{Phase}(a_{\mathrm{ini}})\leq\mathit{Phase}(a_{\mathrm{res}})$
20  if $\mathit{Phase}(a_{\mathrm{res}})=T_{4}$ then
@@
22  Let $p_{1}=\mathit{Phase}(a_{\mathrm{ini}}),\ p_{2}=\mathit{Phase}(a_{\mathrm{res}})$
@@
27
Algorithm 2 $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ at an interaction where initiator $a_{\mathrm{ini}}$ and responder $a_{\mathrm{res}}$ meet.
Table 3: Initialization of each component
component initialization
FindTarget $a.\mathtt{target}=1$, $a.\mathtt{parity}=(T_{1}\bmod{2})$
Detect $a.\mathtt{det}=0$, $a.\mathtt{list}=\emptyset$, $a.\mathtt{susp}=0$
Rank
$a.\mathtt{index}=0$, $a.\mathtt{rank}=\begin{cases}1&\textit{if }a.\mathtt{leader}=1\\ \bot&\textit{otherwise}\end{cases}$,
$a.\mathtt{nonce}=0$, $a.\mathtt{parity}=(T_{3.5}\bmod 2)$
6 Cyclic Execution of Sub-protocols
Using the phase clock [AAE08] described in Section 5.2, $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ invokes the sub-protocols FindTarget, ${\textsc{Detect}}(r,\rho)$, and Rank in phases $T_{1},T_{2}-2]$, $T_{2},T_{3}-2]$, and $T_{3},T_{4}-2]$, respectively (line 24), where
@@ 6 Cyclic Execution of Sub-protocols
and $c_{T}=\Theta(1)$ is a sufficiently large constant to ensure the correctness of each component. If an agent enters phase $T_{4}$, it raises a reset flag (lines 18–19), returning all agents to the initial phase 0 by the epidemic protocol (lines 9–13). Since each phase completes in $\Theta(n\log n)$ interactions w.h.p. by Lemma 6, one cycle of $T_{4}$ phases requires only $\Theta(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p.
A variable $\mathtt{mode}\in\{\bot,\mathsf{F},\mathsf{D},\mathsf{R}\}$ plays a key role in self-stabilization. The modes $\mathsf{F}$, $\mathsf{D}$, and $\mathsf{R}$ correspond to the sub-protocols FindTarget, Detect, and Rank, respectively. When an agent switches its $\mathtt{mode}$ from $x$ to a different mode $y\neq\bot$, it runs the initialization procedure for component $y$. The initialization procedures for Detect, Rank, and FindTarget are described in Sections 4, 7, and 8, respectively. They are also specified in Algorithms 13, and 4 and summarized in Table 3.
Note that ${\textsc{Detect}}(r,\rho)$ requires the target rank $r$; upon entering mode $\mathsf{D}$, each agent $a$ adopts $a.\mathtt{target}$ for $r$. Mode switching occurs in phases $T_{1}-1$, $T_{2}-1$, and $T_{3}-1$, i.e., immediately before each component starts (line 22). An agent $a$ enters mode $\mathsf{R}$ only if $a.\mathtt{det}=1$, i.e., its rank-collision-detected flag is raised, so rank reassignment cannot begin until ${\textsc{Detect}}(r,\rho)$ detects a collision. Each component is executed only when both interacting agents share the same mode; additionally, during mode $\mathsf{D}$ (collision detection), they must have the same target rank (line 23).
The first part of the main function (lines 9–19) brings all agents into phase 0, resetting their modes to $\bot$, within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p., regardless of the initial configuration (Lemma 7). By Lemma 6, this ensures the population completes one full cycle of the three components in the next $\Theta(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p. This reset mechanism relies on two timers, $\mathtt{reset}\in[0,c_{R}\cdot\lceil\lg n\rceil]$ and $\mathtt{delay}\in[0,c_{D}\cdot\lceil\lg n\rceil]$, where a constant $c_{D}=\Theta(1)$ is chosen sufficiently larger than $c_{R}=\Theta(1)$. We omit the detailed explanation of the update rules for $\mathtt{reset}$ and $\mathtt{delay}$ (lines 9–19) in the main text, as they follow standard techniques [BCC+21a, SEIM21], in particular the PropagateReset subprotocol of [BCC+21a]; their precise definitions appear in the pseudocode.
@@
Let $\mathcal{I}$ be the set of configurations in which $a.\mathtt{clock}=a.\mathtt{reset}=a.\mathtt{delay}=a.\mathtt{det}=0$ and $a.\mathtt{mode}=\bot$ hold for all $a\in A$.
@@
Proof.
Follows from Lemma 1 of [ADK+17], Lemma 5 of [SOK+20], and Lemma 6. ∎
@@
For each $X\in\{\mathsf{F},\mathsf{D},\mathsf{R}\}$, let $\mathcal{C}_{\mathrm{init}}(X)$ denote the set of configurations of $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ in which every agent
@@
    satisfies $\mathtt{reset}=0\land(\mathtt{leader}=1\to\mathtt{delay}=0)$,
  2. 2.
@@
  3. 3.
    has all component-specific variables initialized to their specified initial values. (See Table 3 for the specified initial values.)